For the gas phase decomposition of dinitrogen pentoxide, N2O52 NO2 + 1/2 O2 the rate constant in s-1 has been determined at several temperatures. When ln k is plotted against the reciprocal of the Kelvin temperature, the resulting linear plot has a slope of -1.24×104 K and a y-intercept of 31.3. The value of the rate constant for the gas phase decomposition of dinitrogen pentoxide at 322 K is_______ s-1.
Answer #1

We have the Arrhenius equation:

Taking natural logarithm on both sides:

This equation is similar to the straight line equation y = mx + c

Given that:

Slope = -Ea/R = -1.24 × 104 K

Temperature, T = 322 K

Y - intercept, ln A = 31.3

Substituting given values in equation: